OpenStudy (anonymous):

simplify the expression. 3a^2c/-9a^2c^4

OpenStudy (saifoo.khan):

\[\Large \frac{ 3a^2 c}{-9a^2 c^2} \to \frac{ \cancel {a^2}c}{-3\cancel {a^2}c^2 }\]

OpenStudy (saifoo.khan):

Finally, \[\frac{1}{-3c}\]

OpenStudy (anonymous):

thanks! didn't know what to do with that question.

OpenStudy (saifoo.khan):

So you understand it now?

OpenStudy (anonymous):

yes

OpenStudy (saifoo.khan):

Perfect.